# A stochastic reconstruction theorem on rectangular increments with an application to a mixed hyperbolic SPDE

### v2

@@ 2.3 Commuting filtration
We now recall the probabilistic assumptions to formulate our results. Throughout this paper, we fix a complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. Given $m\geq 1$ the set of $m$- integrable random variables $X\colon\Omega\to{\mathbb{R}}$ will be denoted by $L_{m}(\Omega)$ and its norm by $\|\cdot\|_{m}$. We will keep this discussion rather short, as it has been done in length in [BHK23]. One can find more classical references in [Imk85, Wal86, Kho02].
We equip our probability space with a filtration associated with the rectangular increment setting.
@@
The stochastic reconstruction theorem, as the stochastic sewing lemma, is based on BDG-type inequalities which in $d=1$ can be found in [Le20], Equation (2.5). The multiparameter version of this inequality has been found in [BHK23, Lemma 25].
@@ 3.1 Hölder spaces for random distributions
We start by recalling the deterministic multiparameter Hölder spaces, see [QST07, Har21] for further references.
Given ${\bm{\alpha}}\in[0,1]^{d}$ we set $C^{\bm{\alpha}}$ to be the space of all maps $Y\colon[0,T]^{d}\to\mathbb{R}$, such that
@@
To extend these spaces when there is a filtered probability space $(\Omega,\mathcal{F},P)$ with a commuting $d$-parameter filtration $(\mathcal{F}_{\mathbf{t}})_{\mathbf{t}\in[0,T]^{d}}$, we need to condition the random distributions over a careful choice of filtrations. More specifically, we want to quantify the seminorm $\square^{\theta}_{\mathbf{x},\mathbf{y}}Y$ in (3.1) with a $L_{m}$ norm and additionally we want the conditional expectation $\mathbb{E}^{\eta}_{\mathbf{x}}$ to add another regularity factor.
@@
Note that $\left\llbracket Y\right\rrbracket$ is not a norm on $Y$ but only a semi-norm over $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$. For ${\bm{\delta}}={\bm{0}}$, we write $C^{{\bm{\alpha}}}L_{m}:=C^{{\bm{\alpha}},{\bm{0}}}L_{m}$. Further observe that strong martingales as described by [Imk85], [Wal86] fulfill (3.2) uniformly for all ${\bm{\delta}}>{\bm{0}}$. In this case, we set
@@
and write $Y\in C^{{\bm{\alpha}},\infty}L_{m}$ if $\left\|Y\right\|_{C^{{\bm{\alpha}},\infty}L_{m}}<\infty$.
@@
Remark 3.5.
Let us give an easy but useful result regarding the sum of distributions in different $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ spaces.
@@
where ${\bm{\lambda}},\mathbf{x},\psi,\eta$ are chosen as in (3.6). One checks that $\left\|(f+g)(\psi_{\mathbf{x}}^{\bm{\lambda}})\right\|_{m}$ follows a similar bound. Straightforward calculation then gives (3.7). ∎
@@
The most simple examples of elements of $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ spaces are Brownian sheet and white noise. A Brownian sheet is a centered Gaussian process $B_{\mathbf{t}}$, $\mathbf{t}\in\mathbb{R}_{+}^{d}$ characterized by the covariance structure
see for example [Kho02, CW75]. It especially fulfills for any $m\geq 1$
for some constant $c$ depending on $m$ and all $\mathbf{s}\leq\mathbf{t},\theta\subset[d]$. Its conditional expectation fulfills for any $\emptyset\neq\eta\subset\theta$
$\mathbb{E}^{\eta}_{\mathbf{s}}\square^{\theta}_{\mathbf{s},\mathbf{t}}B=0\,.$
This implies $B\in C^{\frac{1}{2}{\bm{1}},\infty}L_{m}$.
Its distributional derivative is given by white noise. More generally, we define white noise as the centered Gaussian process $\xi(\psi)$, $\psi\in L_{2}(\mathbb{R}_{+}^{d})$ with covariance structure
[…]
for any $\psi\in L_{2}(\mathbb{R}^{d}),\mathbf{x}\in\mathbb{R}^{d}_{+}$ and ${\bm{\lambda}}\in(0,1]^{d}$, such that $\psi_{\mathbf{x}}^{\bm{\lambda}}$ has support in +$\mathbb{R}^{d}_{+}$ and some universal constants $c(m)$ depending only on $m$. It follows together with
$\mathbb{E}^{\eta}_{\mathbf{x}}\xi(\psi_{\mathbf{x}}^{\bm{\lambda}})=0\,,$
for $\eta\subset[d]$ and $\mbox{supp}(\psi)\subset[0,1]^{d}$ that $\xi\in C^{-\frac{1}{2}{\bm{1}},\infty}L_{m}$.
@@ 3.2 Coherent stochastic germs and main statement
We now pass to the main statement of the article. Similar to [CZ20, Ker23], we first introduce the notion of coherence, which is the minimal assumption under which reconstruction is possible. For any ${\bm{\alpha}}<{\bm{0}}$, we choose $r>0$ to be the smallest integer, such that $r+\alpha_{i}>0$ holds for all $i\in[d]$. Throughout this section, we assume $(\mathcal{F}_{\mathbf{t}})_{\mathbf{t}\in[0,T]^{d}}$ to be a commuting filtration.
@@
Let ${\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}\in{\mathbb{R}}^{d}$ with ${\bm{\delta}}\geq{\bm{0}}$ and ${\bm{\gamma}}\geq{\bm{\alpha}}$. An $(\mathcal{F}_{\mathbf{x}})_{\mathbf{x}\in[0,T]^{d}}$-adapted germ $(F_{\mathbf{x}})$ is called stochastic $({\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}})$-coherent, if for all $\eta\subset\theta\subset[d]$, $\psi\in B^{r}([0,1]^{\theta}\times[-1,1]^{\theta^{c}})$ and $\mathbf{x},\mathbf{y}\in[0,T]^{d}$ such that $\mathbf{x}_{\theta}\leq\mathbf{y}_{\theta}$ and $\mbox{supp}(\psi_{\mathbf{x}}^{\bm{\lambda}})\in[0,T]^{d}$, it holds that
@@
The following result shows that this is equivalent to our definition:
@@
Note that we switch to test functions with $rd$ derivatives, as we use Lemma 2.8 in the proof of the reconstruction theorem.
@@ 3.3 Proof of the stochastic reconstruction theorem
see [Ker23], pages 19-20 for the details. This procedure only adds a constant depending on $R$ and the dimension $d$. ∎
@@
which shows (3.10). We further note that all $\mathcal{R}^{\kappa}_{\mathbf{x}}(F)$ are linear maps from $C_{c}^{rd}\to\mathbb{R}$ and by the above inequality bounded in $\left\|\psi\right\|_{C_{c}^{rd}}$, so they are random distributions. By construction, $\mathcal{R}^{\kappa}_{\mathbf{x}}(F)$ is adapted and $\mathcal{R}_{\mathbf{x}}^{\kappa}(F)=\mathcal{R}_{\pi_{\mathbf{x}}^{\kappa}{\bm{0}}}(F)$. We finally show
@@
We already know that $\mathcal{R}^{\emptyset,\mathbf{n}}_{\mathbf{x}}$ is a Cauchy-sequence and by [Ker23, Lemma 62], it follows that $F_{\mathbf{x}}=\lim_{n\to\infty}F_{\mathbf{x}}(P_{n{\bm{1}}}\psi)$, so the limit needs to be $F_{\mathbf{x}}$. This finishes the proof of convergence and to show the properties in Theorem 3.12.
@@
which implies that $Z^{\theta}_{\mathbf{x}}(\psi)=\lim_{n\to\infty}Z^{\theta}_{\mathbf{x}}(P_{n{\bm{1}}}\psi)=0$, where we used [Ker23, Lemma 62]. ∎
@@
It is interesting to see that the stochastic parameter ${\bm{\delta}}$ is lost at the level of the reconstruction, i.e. $\mathcal{R}(F)$ will in general not be in $C^{\alpha,{\bm{\delta}}}L_{m}$. This is natural, as $\mathcal{R}(F)$, when evaluated against $\psi^{{\bm{\lambda}}}_{\mathbf{x}}$ locally behaves like $F_{\mathbf{x}}(\psi^{{\bm{\lambda}}}_{\mathbf{x}})$ which is only in $C^{\alpha}L_{m}$. However, the multiplication with a (stochastic) white noise $F=(Y\cdot\xi)$ as in (1.3) does result in $\mathcal{R}(F)\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ (in fact even $\mathcal{R}(F)\in C^{{\bm{\alpha}},\infty}L_{m}$), see Corollary 4.6 for details.
@@ 4 Application to a mixed SPDE
Using the results established in the previous section we will now show the existence and uniqueness of the equation (1.8). As explained in the introduction, we can write (1.8) into its mild formulation to obtain formally
$u(\mathbf{x})=\mathcal{I}(v)(\mathbf{x})+\int_{{\bm{0}}}^{\mathbf{x}}\sigma(u(\mathbf{y}))\xi(d\mathbf{y})+\int_{\bm{0}}^{\mathbf{x}}f(\mathbf{y})u(\mathbf{y})\frac{\partial^{d}}{\partial x_{1}\ldots\partial x_{d}}Z(d\mathbf{y})\,,$
@@
$\mathcal{I}(v)(\mathbf{x})=\sum_{\theta\subset\{1,\ldots d\},\;\theta\neq[d]}(-1)^{1+\sharp(\theta^{c})}v(\pi^{\theta}_{\mathbf{x}}{\bm{0}})\,.$ (4.1)
@@
$u(\mathbf{x})=\mathcal{I}(v)(\mathbf{x})+\int_{{\bm{0}}}^{\mathbf{x}}\mathcal{R}(\sigma(u)\cdot\xi+(fu)\cdot\frac{\partial^{d}}{\partial x_{1}\ldots\partial x_{d}}Z)(d\mathbf{y})\,.$ (4.2)
@@ 4.1 Composition and scalar multiplication with functions
We recall that for any given smooth function $g\colon\mathbb{R}\to\mathbb{R}$ and any $u\in C^{{\bm{\alpha}}}$, $g(u)\in C^{\bm{\alpha}}$, i.e. composition with smooth functions preserves the regularity (For reference, see [QST07, Lemma 3.1] or [BHR22, Lemma 15]). This, however, is no longer the case in the stochastic setting. This is due to the fact, that one would like to apply the Taylor formula to $g(u)$ and use $\square^{\{1\}}_{\mathbf{x},\mathbf{y}}u\square^{\{2\}}_{\mathbf{x},\mathbf{y}}u$ to bound $\square^{(1,2)}_{\mathbf{x},\mathbf{y}}g(u)$. While for deterministic $u$ this gives the correct bounds for $\left|\square^{\{1,2\}}_{\mathbf{x},\mathbf{y}}g(u)\right|$, in the stochastic setting $\square^{\{1\}}_{\mathbf{x},\mathbf{y}}u\square^{\{2\}}_{\mathbf{x},\mathbf{y}}u$ is no longer in $L_{m}$, but only in $L_{m/2}$. Due to this lack of integrability, we can only hope for bounds in $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m/d}$. To keep the same index $m$ invariant we will need to lower the indices ${\bm{\alpha}}$ and ${\bm{\delta}}$. In what follows for $\epsilon\in(0,1]$ we will use the notation $\mathcal{C}^{\epsilon}$ to denote the set of all $\epsilon$-Holder functions $g\colon\mathbb{R}\to\mathbb{R}$ with the convention of taking Lipschitz functions when $\epsilon=1$.

### v3

@@ 2.3 Commuting filtration
We now recall the probabilistic assumptions to formulate our results. Throughout this paper, we fix a complete probability space $(\Omega,\mathcal{F},\mathbb{P})$. Given $m\geq 1$ the set of $m$- integrable random variables $X\colon\Omega\to{\mathbb{R}}$ will be denoted by $L_{m}(\Omega)$ and its norm by $\|\cdot\|_{m}$. We will keep this discussion rather short, as it has been done in length in [BHK23]. One can find more classical references in [Imk85, Wal86, Kho02]. We equip our probability space with a filtration which is compatible with the use of rectangular increments.
@@
The stochastic reconstruction theorem, as the stochastic sewing lemma, is based on BDG-type inequalities which in $d=1$ can be found in [Lê20], Equation (2.5). The multiparameter version of this inequality has been found in [BHK23, Lemma 25].
@@ 3.1 Hölder spaces for random distributions
We start by recalling the deterministic multiparameter Hölder spaces, see [QST07, Har21] for further references. Given ${\bm{\alpha}}\in[0,1]^{d}$ we set $C^{\bm{\alpha}}$ to be the space of all maps $Y\colon[0,T]^{d}\to\mathbb{R}$, such that
@@
To extend these spaces when there is a filtered probability space $(\Omega,\mathcal{F},\mathbb{P})$ with a commuting $d$-parameter filtration $(\mathcal{F}_{\mathbf{t}})_{\mathbf{t}\in[0,T]^{d}}$, we need to condition the random distributions over a careful choice of filtrations. More specifically, we want to quantify the seminorm $\square^{\theta}_{\mathbf{x},\mathbf{y}}Y$ in (3.1) with a $L_{m}$ norm and additionally we want the conditional expectation $\mathbb{E}^{\eta}_{\mathbf{x}}$ to add another regularity factor.
@@
Note that $\left\llbracket Y\right\rrbracket$ is not a norm on $Y$ but only a semi-norm over $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$. For ${\bm{\delta}}={\bm{0}}$, we write $C^{{\bm{\alpha}}}L_{m}:=C^{{\bm{\alpha}},{\bm{0}}}L_{m}$.
The most simple examples of element of $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ spaces is a Brownian sheets $B=\{B_{\mathbf{t}}\colon\mathbf{t}\in\mathbb{R}_{+}^{d}\}$ together with its canonical filtration. By definition, a Brownian sheet is a centred Gaussian process characterized by the covariance structure
see for example [Kho02, CW75]. It especially fulfills for any $m\geq 1$
for some constant $c(m)\geq 0$ depending on $m$ and all $\mathbf{s}\leq\mathbf{t},\theta\subset[d]$. Considering its conditional expectation with respect to the canonical filtration of the process, for any $\emptyset\neq\eta\subset\theta$ one has
$\mathbb{E}^{\eta}_{\mathbf{s}}\square^{\theta}_{\mathbf{s},\mathbf{t}}B=0\,.$
This condition is due to the property of being a strong martingales, as described by [Imk85, Wal86]. To take in account the “infinite regulation” of these elements, we set
@@
and write $Y\in C^{{\bm{\alpha}},\infty}L_{m}$ if $\left\|Y\right\|_{C^{{\bm{\alpha}},\infty}L_{m}}<\infty$, so that one has $B\in C^{\frac{1}{2}{\bm{1}},\infty}L_{m}$. Another simple example of an element in $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ is given by any random field of the form $Y_{\mathbf{t}}=B_{\mathbf{t}}+Z_{\mathbf{t}}$ with $Z\in C^{\bm{\beta}}$ for $\beta\geq 1/2{\bm{1}}$. In this case one has $Y\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ with ${\bm{\delta}}=\beta-1/2{\bm{1}}$.
@@
Similarly to the case when ${\bm{\alpha}}>0$, in this setting we can simple explain the elements of $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ with negative ${\bm{\alpha}}$ by simply applying a mixed distributional derivative of the previous examples. In case of Brownian sheet over $\mathbb{R}_{+}^{d}$ we obtain white noise over $\mathbb{R}_{+}^{d}$ that is a the centered Gaussian process $\xi(\psi)$, $\psi\in L_{2}(\mathbb{R}_{+}^{d})$ with covariance structure given by
[…]
for any $\psi\in L_{2}(\mathbb{R}^{d}),\mathbf{x}\in\mathbb{R}^{d}_{+}$ and ${\bm{\lambda}}\in(0,1]^{d}$, such that $\psi_{\mathbf{x}}^{\bm{\lambda}}$ has support in +$\mathbb{R}^{d}_{+}$. It follows also that
$\mathbb{E}^{\eta}_{\mathbf{x}}\xi(\psi_{\mathbf{x}}^{\bm{\lambda}})=0\,,$
for $\eta\subset[d]$ and $\mbox{supp}(\psi)\subset[0,1]^{d}$, thereby obtaining $\xi\in C^{-\frac{1}{2}{\bm{1}},\infty}L_{m}$. Simple examples of elements in $C^{-\frac{1}{2}{\bm{1}},{\bm{\delta}}}L_{m}$ with finite ${\bm{\delta}}$ are given by random distribution of the form $\xi+\zeta$ where $\zeta\in C^{{\bm{\beta}}}$ with $-1/2{\bm{1}}<{\bm{\beta}}<{\bm{0}}$. We conclude by providing a useful result regarding the sum of distributions in different $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$ spaces.
@@
where ${\bm{\lambda}},\mathbf{x},\psi,\eta$ are chosen as in (3.6). One checks that $\left\|(f+g)(\psi_{\mathbf{x}}^{\bm{\lambda}})\right\|_{m}$ follows a similar bound. A straightforward calculation then gives (3.7). ∎
@@ 3.2 Coherent stochastic germs and main statement
We now pass to the main statement of the article. Similar to [CZ20, Ker24], we first introduce the notion of coherence, which is the minimal assumption under which reconstruction is possible. For any ${\bm{\alpha}}<{\bm{0}}$, we choose $r>0$ to be the smallest integer, such that $r+\alpha_{i}>0$ holds for all $i\in[d]$. Throughout this section, we assume $(\mathcal{F}_{\mathbf{t}})_{\mathbf{t}\in[0,T]^{d}}$ to be a commuting filtration.
@@
Let ${\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}}\in{\mathbb{R}}^{d}$ with ${\bm{\delta}}\geq{\bm{0}}$ and ${\bm{\gamma}}\geq{\bm{\alpha}}$. An $(\mathcal{F}_{\mathbf{x}})_{\mathbf{x}\in[0,T]^{d}}$-adapted germ $(F_{\mathbf{x}})$ is called stochastic $({\bm{\alpha}},{\bm{\gamma}},{\bm{\delta}})$-coherent, if for all $\eta\subset\theta\subset[d]$, $\psi\in B^{r}([0,1]^{\theta}\times[-1,1]^{\theta^{c}})$ and $\mathbf{x},\mathbf{y}\in[0,T]^{d}$ such that $\mathbf{x}_{\theta}\leq\mathbf{y}_{\theta}$ and $\mbox{supp}(\psi_{\mathbf{y}}^{\bm{\lambda}})\in[0,T]^{d}$, it holds that
@@
The following result shows some equivalent ways to state the coherence.
@@
Note that we switch to test functions with $rd$ derivatives, because we want to use Lemma 2.8 in the proof of the reconstruction theorem.
@@ 3.3 Proof of the stochastic reconstruction theorem
see [Ker24], pages 19-20 for the details. This procedure only adds a constant depending on $R$ and the dimension $d$. ∎
@@
which shows (3.10). We further note that all $\mathcal{R}^{\kappa}_{\mathbf{x}}(F)$ are linear maps from $C_{c}^{rd}\to\mathbb{R}$ and by the above inequality bounded in $\left\|\psi\right\|_{C_{c}^{rd}}$, so they are random distributions. By construction, $\mathcal{R}^{\kappa}_{\mathbf{x}}(F)$ is adapted and $\mathcal{R}_{\mathbf{x}}^{\kappa}(F)=\mathcal{R}_{\pi_{\mathbf{x}}^{\kappa}{\bm{0}}}(F)$. We finally show that
@@
We already know that $\mathcal{R}^{\emptyset,\mathbf{n}}_{\mathbf{x}}$ is a Cauchy-sequence and by [Ker24, Lemma 62], it follows that $F_{\mathbf{x}}=\lim_{n\to\infty}F_{\mathbf{x}}(P_{n{\bm{1}}}\psi)$, so the limit needs to be $F_{\mathbf{x}}$. This finishes the proof of convergence and to show the properties in Theorem 3.12.
@@
which implies that $Z^{\theta}_{\mathbf{x}}(\psi)=\lim_{n\to\infty}Z^{\theta}_{\mathbf{x}}(P_{n{\bm{1}}}\psi)=0$, where we used [Ker24, Lemma 62]. ∎
@@
It is interesting to see that the stochastic parameter ${\bm{\delta}}$ is lost at the level of the reconstruction, i.e. $\mathcal{R}(F)$ will in general not be in $C^{\alpha,{\bm{\delta}}}L_{m}$. This is natural, as $\mathcal{R}(F)$, when evaluated against $\psi^{{\bm{\lambda}}}_{\mathbf{x}}$ locally behaves like $F_{\mathbf{x}}(\psi^{{\bm{\lambda}}}_{\mathbf{x}})$ which is only in $C^{\alpha}L_{m}$. However, the multiplication with a (stochastic) white noise $F=(Y\cdot\xi)$ as in (1.3) does result in $\mathcal{R}(F)\in C^{{\bm{\alpha}},{\bm{\delta}}}L_{m}$, see Corollary 4.6 for details.
@@ 4 Application to a mixed SPDE
Using the results established in the previous section we will now show the existence and uniqueness of the equation (1.7). As explained in the introduction, we can write (1.7) into its integral form to obtain formally
$u(\mathbf{x})=I(v)(\mathbf{x})+\int_{{\bm{0}}}^{\mathbf{x}}\sigma(u(\mathbf{y}))\xi(d\mathbf{y})+\int_{\bm{0}}^{\mathbf{x}}f(\mathbf{y})u(\mathbf{y})\frac{\partial^{d}}{\partial x_{1}\ldots\partial x_{d}}Z(d\mathbf{y})\,,$
@@
$I(v)(\mathbf{x})=\sum_{\theta\subset\{1,\ldots d\},\;\theta\neq[d]}(-1)^{1+\sharp(\theta^{c})}v(\pi^{\theta}_{\mathbf{x}}{\bm{0}})\,.$ (4.1)
@@
$u(\mathbf{x})=I(v)(\mathbf{x})+\int_{{\bm{0}}}^{\mathbf{x}}\mathcal{R}(\sigma(u)\cdot\xi+(fu)\cdot\frac{\partial^{d}}{\partial x_{1}\ldots\partial x_{d}}Z)(d\mathbf{y})\,.$ (4.2)
@@ 4.1 Composition and scalar multiplication with functions
We recall that for any given smooth function $g\colon\mathbb{R}\to\mathbb{R}$ and any $u\in C^{{\bm{\alpha}}}$, $g(u)\in C^{\bm{\alpha}}$, i.e. composition with smooth functions preserves the regularity (For reference, see [QST07, Lemma 3.1] or [BHR24, Lemma 15]). This, however, is no longer the case in the stochastic setting. This is due to the fact, that one would like to apply the Taylor formula to $g(u)$ and use $\square^{\{1\}}_{\mathbf{x},\mathbf{y}}u\square^{\{2\}}_{\mathbf{x},\mathbf{y}}u$ to bound $\square^{(1,2)}_{\mathbf{x},\mathbf{y}}g(u)$. While for deterministic $u$ this gives the correct bounds for $\left|\square^{\{1,2\}}_{\mathbf{x},\mathbf{y}}g(u)\right|$, in the stochastic setting $\square^{\{1\}}_{\mathbf{x},\mathbf{y}}u\square^{\{2\}}_{\mathbf{x},\mathbf{y}}u$ is no longer in $L_{m}$, but only in $L_{m/2}$. Due to this lack of integrability, we can only hope for bounds in $C^{{\bm{\alpha}},{\bm{\delta}}}L_{m/d}$. To keep the same index $m$ invariant we will need to lower the indices ${\bm{\alpha}}$ and ${\bm{\delta}}$. In what follows for $\epsilon\in(0,1]$ we will use the notation $\mathcal{C}^{\epsilon}$ to denote the set of all $\epsilon$-Holder functions $g\colon\mathbb{R}\to\mathbb{R}$ with the convention of taking Lipschitz functions when $\epsilon=1$.